### v1

# Robust Understanding of Human-Robot Social Interactions through Multimodal Distillation

Abstract.
The need for social robots and agents to interact and assist humans is growing steadily. To be able to successfully interact with humans, they need to understand and analyse socially interactive scenes from their (robot’s) perspective. Works that model social situations between humans and agents are few; and even those existing ones are often too computationally intensive to be suitable for deployment in real time or on real world scenarios with limited available information. We propose a robust knowledge distillation framework that models social interactions through various multimodal cues, yet is robust against incomplete and noisy information during inference. Our teacher model is trained with multimodal input (body, face and hand gestures, gaze, raw images) that transfers knowledge to a student model that relies solely on body pose. Extensive experiments on two publicly available human-robot interaction datasets demonstrate that the our student model achieves an average accuracy gain of 14.75% over relevant baselines on multiple downstream social understanding task even with up to 51% of its input being corrupted. The student model is highly efficient: it is $<1$% in size of the teacher model in terms of parameters and uses $\sim 0.5$‰ FLOPs of that in the teacher model. Our code will be made public during publication.
Human-Robot Interaction, Multimodal Learning, Knowledge Distillation, Egocentric Vision
ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing HCI theory, concepts and models
Figure 1. Our aim is to build a lightweight and robust model to understand human-robot social interactions. A novel teacher-student learning framework is built to achieve this, where the teacher model uses multimodal cues (body pose, face and hand gestures, and gaze) and raw RGB frames to learn a rich representation of the social scenario. Through knowledge distillation, the teacher model is able to transfers the knowledge to a lightweight student model that relies solely on body pose features. The student model is robust against incomplete and noisy inputs.
1. Introduction
Recent application advances in social robotics and virtual agents include personal care, navigation, and user interaction (Zhou et al., 2022). A central challenge in social AI and human-robot interaction (HRI) is enabling the systems to understand and forecast human social behaviours from an egocentric perspective, such as intent to interact, attitudes, and social actions (Bian et al., 2025). This capability is particularly critical during the early stages of interaction, where understanding the user’s engagement intent allows the system to proactively prepare for smooth and natural interactions. Early prediction enhances user experience and supports adaptive, context-aware responses (Avelino et al., 2021). However, deploying such predictive models in real-world settings remains challenging due to constraints in computational resources, input quality, and the need for real-time processing.
Understanding human social behaviours relies on multiple non-verbal signals (Urakami and Seaborn, 2023), including body, face and hand gestures, gaze and raw images, each providing unique and complementary information into social interactions (Yumak and Magnenat-Thalmann, 2015). Despite progress in understanding multimodal nonverbal cues, several challenges remain in using them for broader social understanding in real-world deployment. (1) Multimodal Integration: Social understanding needs fusing information from diverse modalities that can sometimes be inconsistent. Models need to integrate information and differences across modalities to extract coherent semantic representations. (2) Real-World Challenges: Egocentric perspectives introduce challenges such as limited field of view, camera motion, occlusions, and feature estimation errors, which degrades the reliability of social understanding. (3) Computational Efficiency: Using multiple modalities requires computationally intensive models and more feature estimation modules, posing difficulties for real-time inference on resource-constrained systems, such as social robots. Addressing these challenges needs a lightweight yet robust framework that can integrate multimodal cues under noisy conditions while supporting real-time social understanding.
To tackle these challenges, we propose a knowledge distillation framework that enables a lightweight student model to infer rich social cues from incomplete body pose input alone and jointly forecast three key aspects of social understanding: interaction intentions, attitudes, and social actions, as shown in Fig.1. Our approach consists of two stages: first, we train a teacher model with full-modal inputs, including body, face and hand gestures, gaze and raw images, to learn and output high-quality social representations. Then, we distil this knowledge into a lightweight student model that relies solely on body pose while incorporating spatio-temporal noise to simulate real-world data degradation. Instead of only learning from downstream task labels, the student also learns to infer multimodal and clean social cues through knowledge distillation, ensuring robust performance under noisy conditions. The lightweight design of the student model improves computational efficiency.
In summary, our work enhances the robustness and efficiency of jointly forecasting users’ intent to interact, attitudes, and social actions from an egocentric perspective through the contributions:
  • 1
    Knowledge distillation framework: We propose a knowledge distillation framework where a teacher model SocialC3D using body, face and hand gestures, gaze and raw images, supervises a lightweight student model SocialEgoMobile that only uses body pose features. The average accuracy of the student model improved by 9.13% over independent training on JPL-Social (Bian et al., 2025) and 11.32% on the HARPER (Avogaro et al., 2024).
  • 2
    Robustness to real-world noises: To enhance robustness against noisy and incomplete input, we introduce spatiotemporal corruption for the student model. With knowledge distillation, when 51% of spatiotemporal information is lost the student maintains an average accuracy of 81.35% on JPL-Social (Bian et al., 2025) (14.68% improvement over independent training) and 59.05% on HARPER (Avogaro et al., 2024) (14.82% improvement).
  • 3
    Efficiency and deployment readiness: The proposed student model SocialEgoMobile demonstrates high computational efficiency, requiring only 1% of model parameters and 0̃.5‰ of the floating-point operations (FLOPs) compared to the teacher model.
2. Related Work
2.1. Social Intent Prediction
Egocentric social intent prediction is crucial for socially intelligent agents to establish natural, harmonious, and efficient interactions with users. A precise understanding of the user’s social behaviour enables agents to timely adjust response strategies to meet user needs. Current research primarily models social intent as a binary classification task, focusing on predicting whether a user intends to interact with an intelligent agent. Approaches include analyzing motor cues such as body orientation, movement direction, and speed (Abbate et al., 2024), targeting users interested in interaction using visual and audible features (Pourmehr et al., 2017), and using gaze information in virtual environments to infer interaction intent (David-John et al., 2021). However, these methods are limited to distinguishing between ‘interact’ and ‘not interact’, failing to capture the richness and complexity of users’ social intentions.
Human social intent estimation is inherently more nuanced than binary classifications and includes wider aspects. Recent research has explored various dimensions of users’ social behaviours. Convolutional neural networks (CNNs) have been used to analyze facial expressions and body pose for attitude and emotion perception (Ilyas et al., 2021), while multilayer perceptions (MLPs) have been utilized to process facial and thermal images to infer user emotions during HRI (Filippini et al., 2021). Gated recurrent units (GRUs) capture the temporal dependencies of facial, acoustic, and textual features, enabling the detection of dynamic changes in users’ attitudes and emotions during conversations (Li et al., 2020). Similarly, interaction action prediction has gained attention. Using cascade histogram of time series gradients can capture past spatio-temporal features and predict users’ future social actions (Ryoo et al., 2015). Social action and wider social signal in multi-agent scenarios have also been explored, by integrating body, facial, and hand pose features to capture user’s body movement (Bian et al., 2025).
2.2. Non-Verbal Cues in Multimodal HRI
In Human-Robot Interaction (HRI), accurately interpreting the user’s identity, social intent, behaviours, and communication requires integrating information from multiple modalities (Yumak and Magnenat-Thalmann, 2015). Meanwhile, non-verbal cues play a crucial role in understanding user intentions and needs (Urakami and Seaborn, 2023). For instance, while detecting sarcasm through text is challenging during a conversation, recognizing subtle body movements such as shoulder shrugs or lip curls allows humans to accurately identify sarcasm (Castro et al., 2019). Consequently, capturing and interpreting mutlimodal non-verbal cues in HRI is essential for assessing the users. Recent research has explored various strategies for multi-modal fusion to enhance the accuracy of interaction models. For example, combining RGB images, facial expressions, and distance information can effectively assess user engagement (Ben-Youssef et al., 2019) and interest (Foster et al., 2019) when interacting with service robots in public spaces. Facial expression and gaze information can be used to infer whether a user requires assistance in collaborative tasks (Wilson et al., 2022). To further improve robustness in real-world scenarios, researchers tried to model temporal information of non-verbal signals. Recurrent models have been used to model gestures and RGB video streams for identifying user action commands (Liu et al., 2018), and to detect user emotions during collaboration by analyzing body movements and facial expressions (Páez and González, 2022). Pose skeletons and RGB information have been fused to predict interaction actions in single-user (Islam and Iqbal, 2020) and multi-user scenarios (Yasar et al., 2024). Non-verbal modalities including body, face, hand gestures, gaze and raw images provide rich contextual information that enhances a robot’s ability to accurately understand to users’ intentions, attitudes, and actions, improving the effectiveness of HRI systems.
Research gap. Despite the progress in social understanding using deep learning models, existing approaches often rely on multimodal inputs. However, processing more modal information increases computational cost (Liang et al., 2024) and the deep models are sensitive to missing or noisy information (Drenkow et al., 2021), making them impractical for deployment on resource-constrained devices and real-world scenarios. While knowledge distillation has been explored to compress large models into lighter versions, most prior works focus on preserving accuracy under ideal conditions, overlooking the need to integrate multimodal information using a single modality and maintain robustness in environments with incomplete or degraded inputs (Arani et al., 2021). To address these limitations, we propose a knowledge distillation framework with a lightweight, body pose-only student model that learns from a full-modality teacher model through a robust distillation, ensuring both computational efficiency and resilience to temporal and spatial noise.
3. Methodology
Figure 2. Our knowledge distillation framework uses SocialC3D as the teacher model, which fuses raw images, body, face, hand gestures, and gaze. Each modality is processed by a ResNet (He et al., 2016) and integrated via lateral connections and late fusion, producing a high-quality social representation for downstream tasks. The lightweight student model, SocialEgoMobile, uses only corrupted body pose. It consists of a two-layer GAT (Veličković et al., 2018) and a single-layer Bi-LSTM (Graves and Schmidhuber, 2005) to extract social representations. The framework distillates knowledge from the teacher model by maximising the mutual information of the social representations output by the teacher and student model.
3.1. Problem Definition
Given a sequence of egocentric videos captured by an intelligent social agent, our aim is to forecast the user’s social intent and behaviours in real time based only on the early stages of interaction (e.g., the first second). Specifically, we forecast the user’s intent to interact, attitude, and social actions. To achieve this, we propose a knowledge distillation framework where a lightweight student model is trained to rely on body pose features, while a teacher model uses full multimodal inputs (body, face, hand gesture, gaze and raw images), as shown in Fig.2. Body pose features, facial keypoints, and hand keypoints are extracted from the video using Alphapose (Fang et al., 2022) and gaze features are extracted using MCGaze (Guan et al., 2023). Formally, we define the input signals at step $t$ as follows:
  • $\boldsymbol{x}_{t}^{\text{Pose}}$: The coordinates of 17 body pose keypoints
  • $\boldsymbol{x}_{t}^{\text{Image}}$: The raw images
  • $\boldsymbol{x}_{t}^{\text{Face}}$: The coordinates of 68 facial keypoints
  • $\boldsymbol{x}_{t}^{\text{Hand}}$: The coordinates of 42 hands keypoints
  • $\boldsymbol{x}_{t}^{\text{Gaze}}$: The coordinates of the beginning and end of gaze
The teacher model $F_{teacher}$ is trained with the complete set of input modalities:
(1) $R_{teacher}=F_{teacher}(\boldsymbol{x}_{t:T}^{\text{Pose}},\boldsymbol{x}_{t:T}^{\text{Image}},\boldsymbol{x}_{t:T}^{\text{Face}},\boldsymbol{x}_{t:T}^{\text{Hand}},\boldsymbol{x}_{t:T}^{\text{Gaze}})$
where $R_{t}$ represents the high-level social representation extracted by the teacher model. To ensure efficiency and robustness, we introduce a lightweight student model $F_{student}$, which relies solely on body pose features $\boldsymbol{x}_{t}^{\text{Pose}}$:
(2) $R_{student}=F_{student}(\boldsymbol{x}_{t:T}^{\text{Pose}})$
where $R_{student}$ represents the social representation extracted by the student model.
For the downstream tasks of forecasting user’s intent to interact, attitude, and social actions, the social representations $R$ extracted by the teacher or student model are processed by the Chain hierarchical classifier (Bian et al., 2025), which simulates how human understand and forecast social signals, predicting the results of three downstrem tasks jointly:
(3) $\mathcal{L}_{intent},\mathcal{L}_{attitude},\mathcal{L}_{action}=Classifier(R)$
3.2. Teacher Model
To effectively integrate multiple modalities, we employ a teacher model, SocialC3D which is based on PoseC3D (Duan et al., 2022), a spatio-temporal feature extraction framework using 3D convolutional neural networks (3D-CNN). PoseC3D converts sequences of 2D pose keypoints into 3D heatmap stacks, according to their coordinates and confidence scores. The heatmaps are converted as follows:
(4) $H_{i}(x,y)=\exp\left(-\frac{(x-x_{i})^{2}+(y-y_{i})^{2}}{2\sigma^{2}}\right)*c_{i}$
where $\sigma$ is the Gaussian variance of the heat map; ($x$, $y$) are the coordinates of the heatmap; ($x_{i}$, $y_{i}$) and $c_{i}$ are the coordinates and confidence scores of the current keypoint $i$. This approach effectively captures spatio-temporal dependencies and exhibits greater robustness to noise than traditional skeleton-based methods. Additionally, the lateral connections between the RGB and Pose pathway fuse image and pose information, enhancing multimodal learning.
Based on this, we introduce SocialC3D to incorporate more modal information. SocialC3D transforms additional modalities, facial expressions, hand gestures, and gaze, into 3D heatmap stacks as input, which are processed using the same ResNet (He et al., 2016) as the body pose stream. For gaze estimation, we use the centre of the head bounding box as the beginning and calculate the end based on the gaze direction and the diagonal length of head bounding box. SocialC3D incorporate lateral connections to propagate intermediate representations across modalities at multiple stages, enabling deeper feature integration.
3.3. Student Model
The student model takes only body pose keypoints as input. Compared to raw visual information, body language provides a more direct and interpretable signal for understanding social interactions. Additionally, body posture inherently contain some facial and hand dynamics, allowing the model to integrate multimodal cues through knowledge distillation. Moreover, body pose features, as high-level abstractions, support lighter spatio-temporal analysis models and pose estimation, enhancing the student model’s suitability for resource-constrained settings.
To reduce computational complexity while maintaining strong predictive performance, we developed SocialEgoMobile as our student model. To reduce the risk of overly deep networks amplifying noise and hindering useful information extraction, we designed a streamlined architecture consisting of a two-layer graphical attention network (GAT) (Veličković et al., 2018) and a single-layer bi-directional long- and short-term memory (Bi-LSTM) (Graves and Schmidhuber, 2005). By dynamically adjusting the attentional weights between neighbouring joints, GAT can effectively capture the spatial relationships between key points of the body pose skeleton graph, thus making the model more focused on useful information and limiting the noise. Meanwhile, Bi-LSTM enhances temporal modelling by processing sequences forward and backward with forgetting gates that can selectively filter noise. This lightweight yet effective design maintains stable performance under noisy and incomplete input conditions, thus enhancing robustness and making it suitable for real-time deployment in resource-limited environments.
3.4. Knowledge Distillation
To effectively transfer knowledge of the user’s social behaviour from the teacher model to the student model, we adopt a feature-based knowledge distillation method. The student model learns not only from the ground truth labels of the downstream tasks but also by minimizing the discrepancy between its high-level social representations and those of the teacher model. We employ InfoNCE (Information Noise Contrastive Estimation) (Oord et al., 2018) to estimate and maximize the mutual information between teachers’ and students’ social representations. Specifically, InfoNCE enhances similarity between social representation of teacher and student models of the same input while reducing similarity between different inputs. During training, the loss function of the student model combines the mutual information loss from InfoNCE with the classification loss from the downstream tasks, weighted for back propagation. The loss of InforNCE is defined as:
(5) ${{Loss}_{\text{infoNCE}}}$ $=-\mathop{{}\mathbb{E}}\left[\log\frac{\exp(f(x,x^{+}))}{\sum_{x_{j}\in X}\exp(f(x,x_{j}))}\right]$
where $x$ is the teacher model’s social representation for a given input, $x^{+}$ is the representation from the student model for the same input, and $x_{j}$ is the representation from the student model for different inputs. The similarity function $f(x,x^{\prime})$ is defined using cosine similarity as follows:
(6) $\mathbf{}{f}(\mathbf{x},\mathbf{x^{\prime}})=\frac{\mathbf{x}\cdot\mathbf{x^{\prime}}}{\|\mathbf{x}\|\|\mathbf{x^{\prime}}\|}$
By knowledge distillation, the student model can still learn high-level social knowledge from the teacher model with limited modality and corrupted inputs. Furthermore, we introduce spatio-temporal corruption in the student model’s input, enhancing its ability to integrate multimodal social cues while maintaining robustness to noise and missing data.
3.5. Input Corruption
In real-world applications, the social understanding model must not only work under the computational constraints of mobile platforms but also remain robust to incomplete and noisy inputs. To simulate such conditions, we deliberately introduce corruption in both spatial and temporal dimensions during training. In the spatial domain, a fixed number of body pose keypoints are randomly selected and corrupted by either setting their coordinate values and confidence scores to zero or replacing the coordinates with random values within the normalized range. This mimics localized body occlusions or pose estimation errors. In the temporal domain, we randomly drop a fixed number of frames by setting all keypoint values in those frames to zero or normalized random noise, simulating transient global occlusions. These corruptions force the student model to learn robust representations and maintain performance under incomplete or noisy input during knowledge distillation.
4. Experiments
4.1. Data
Figure 3. We standardized the input frame rate of two datasets to 10 FPS and limited the observation window to the first second of social interaction.
To evaluate the performance and robustness of our models in forecasting users’ social intents and behaviours from an egocentric perspective, we use two datasets of human-robot social interactions recorded from the robot’s perspective. Examples of datasets are shown in Fig. 3.
JPL-Social (Bian et al., 2025) is an extension of the JPL-Interaction datasets (Ryoo and Matthies, 2013; Ryoo et al., 2015), designed for human-robot interaction research, with a focus on users’ social actions within the central visual field. The original dataset captures 8 social actions performed by the main user: handshake, hug, pet, wave, punch, throw, point, and leave. JPL-Social extends this by annotating additional users in the field of view during multi-person scenarios, introducing two additional social actions: gaze and no response. Furthermore, JPL-Social provides labels for social intent (categorized as interested, interacting, or not interested) and attitude (positive or negative) for each user. The dataset comprises 290 interaction sequences involving 16 participants in diverse environmental settings.
HARPER (Avogaro et al., 2024) is a dataset capturing 10 human-robot social actions performed by 17 users, recorded from multiple perspectives using five cameras mounted on a quadrupedal robot. For our research, we use 6 social actions specifically relevant to robot-oriented interactions: crash, walk-avoid, touch, walk-stop, punch, and kick, so there are 260 interaction sequences. In addition, HARPER includes variations in viewing angle and resolution, increasing the difficulty of understanding users’ social signals. To ensure consistency in social representation, we manually annotated HARPER with the same social intent and attitude labels as JPL-Social.
4.2. Experiment Setting
We set the Frames Per Second (FPS) of the videos to 10 and limited the observation window to the first second of social interaction, which starts when the user approaches the social agent, so the inputs have 10 frames. To expand the training set, we randomly cropped video frames and re-extracted pose features, facial expressions, hand gestures, and gaze information. After horizontal flipping, we augmented number of videos 20 times for richer training data. We train and test the models on the two datasets separately. Given our focus on model efficiency, we report both the number of model parameters and the number of floating point operations (FLOPs) for each input.
4.3. Performance of the Teacher Model
Model Params FLOPs JPL-Social HARPER
(M) (M) Intent Acc. Attitude Acc. Action Acc. Intent Acc. Attitude Acc. Action Acc.
ST-GCN+ (Yan et al., 2018) 43.86 1924.87 86.90 76.19 71.43 86.54 73.08 78.85
ST-TR+ (Plizzari et al., 2021) 58.48 4724.63 79.76 59.52 48.81 75.00 65.38 59.61
MS-G3D+ (Liu et al., 2020) 48.82 5241.97 88.10 80.95 76.19 90.38 78.85 80.77
SocialEgoNet+ (Bian et al., 2025) 37.78 422.35 86.90 77.38 71.43 86.54 75.00 78.85
SocialC3D (ours) 48.49 7902.19 92.85 88.10 82.14 96.15 82.69 88.46
SocialEgoMobile (ours) 0.43 4.23 82.14 71.43 67.86 69.23 44.23 51.92
Table 1. Comparison of SocialC3D and SocialEgoMobile with state-of-the-art methods on the JPL-Social and HARPER. SocialEgoMobile relies solely on clean body pose features as input. ’+’ indicates that the model uses raw image and gaze information.
As the teacher model, we compare SocialC3D with several state-of-the-art methods. Although originally designed for body pose skeleton graphs, these methods can also process facial and hand keypoint graphs. To ensure a fair comparison, we incorporated the raw image and gaze information channels from SocialC3D into all teacher models, aligning the available modalities across teacher models. Each modality channel in the teacher models is trained independently first and then during joint fine-tuning, outputs of all the channels are fused before being passed to the chain hierarchical classifier. To enhance body language understanding, the body pose and raw image channels of all teacher models are initialized with pre-trained weights on Kinetics 400 (Kay et al., 2017).
As shown in Tab le1, we compare SocialC3D with SOTA methods on the JPL-Social and HARPER datasets. SocialC3D consistently achieves superior performance on three downstream social tasks in both datasets. Compared with the second-best model, MS-G3D (Liu et al., 2020), SocialC3D achieves an average accuracy improvement of 5.85% on JPL-Social and 5.77% on HARPER, which highlights the effectiveness of SocialC3D in understanding social cues and forecasting users’ social intentions and behaviours from an egocentric perspective. We attribute this improvement to two key factors. First, instead of directly using pose features, converting them into heatmaps enhances robustness to pose estimation errors, which is particularly pronounced in egocentric settings due to narrow and jittery fields of view. Second, our experiments focus on the initial 10 frames of social interactions. 3D CNNs are suitable for capturing local temporal dynamics, making them more appropriate for our experiment setting. In contrast, models like ST-GCN (Yan et al., 2018) and MS-G3D (Liu et al., 2020) rely on complete and accurate pose features, making them more vulnerable to missing joints and estimation noise. ST-TR (Plizzari et al., 2021), while effective in modelling long-term dependencies due to using transformers (Vaswani et al., 2017), struggles with short sequences and lacks mechanisms to handle input degradation. Although SocialEgoNet (Bian et al., 2025) is designed to capture users’ non-verbal social cues from an egocentric perspective, its lightweight architecture limits its capacity to extract fine-grained information independently.
In addition, we report the performance of SocialEgoMobile using only body pose to compare its accuracy and computational efficiency against the teacher models. Compared to SocialC3D, SocialEgoMobile’s average accuracy of downstream tasks on JPL-Social was 13. 89% lower and 33.97% lower on HARPER. The larger performance gap on HARPER is attributed to its recording setup, where the robot dog has a lower vantage point and fewer frames with full-body visibility, resulting in noisier and more incomplete pose inputs, as shown in Fig.3. In terms of computational efficiency, compared to SocialC3D, SocialEgoMobile only needs 0.98% of the model parameters and 0.54‰of FLOPs. These results show SocialEgoMobile is computationally lightweight enough for social understanding on mobile platforms with limited resources.
4.4. Robustness Analysis on the Student Model
Figure 4. Knowledge distillation (KD) consistently improves the performance of the student model, SocialEgoMobile, under Individual and simultaneous spatio-temporal corruption on all three downstream tasks, interaction intent, attitude, and social action forecast. Improvements on downstream task accuracy through distillation are labelled.
To address the practical challenges faced by mobile socially intelligent agents from an egocentric perspective, such as occlusion of user body parts and pose estimation error, we deliberately introduced random spatio-temporal corruption into the input of the student model SocialEgoMobile, as mention in Section 3.5. The goal is to improve robustness by using the rich and clean knowledge extracted from the intermediate social representations of the teacher model SocialC3D, to supervise the learning of the student model with corrupted inputs. During the robustness tests, student models were trained for 70 epochs. The first 30 epochs are warm-up phase, where the corruption rate was gradually increased to the final value to prevent unstable training due to excessive information loss.
Fig.4 compares SocialEgoMobile’s performance under two settings: independent learning and knowledge distillation. Under independent learning, performance degrades as corruption increases. When 30% of joints and frames is corrupted, $30\%\oplus 30\%=51\%$ of the information in the input was corrupted. The average accuracy across three downstream tasks drops by 8.85%, demonstrating the increased difficulty and robustness demands. In contrast, with knowledge distillation, SocialEgoMobile consistently outperforms the independently trained counterpart across all input settings. When 30% of joints and frames is corrupted, SocialEgoMobile trained via distillation surpasses the performance of the independently trained model using clean input. This suggests that knowledge distillation effectively transfers social knowledge from the teacher model, which incorporates additional modalities and complete, clean inputs, to the student model. So it can more accurately capture valuable social cues even when facing corrupted inputs. The performance gains are particularly significant for the user attitude and social action prediction tasks, which rely more on detailed and accurate input.
Whether trained independently or via knowledge distillation, SocialEgoMobile is more sensitive to the corruption of spatial information than temporal information. This is because temporal patterns in human behaviour tend to be more regular and predictable, allowing models to more effectively compensate for missing frames and filter out noise. In contrast, spatial patterns of pose graphs are more complex, such as body structure and human kinematics.
4.5. Ablation Experiments
Figure 5. Performance comparison on JPL-Social and HARPER of SocialC3D using different input modalities, and the impact of corresponding SocialC3D-based Knowledge Distillation on SocialEgoMobile (SEM) when 30% of spatio-temporal information are corrupted. Accuracy drops due to missing modalities are labelled.
Contribution of Mutli-Modalities.  We incorporate body pose, raw images, face, hand gestures, and gaze as input modalities for the teacher model SocialC3D, enabling it to comprehensively capture social cues from users’ non-verbal behaviours across multiple perspectives. Body pose and raw images are the foundational modalities and prior work has demonstrated their effectiveness and complementarity in understanding body movements (Duan et al., 2022). In contrast, facial expressions, hand gestures, and gaze focus on localized regions of the body, providing fine-grained cues essential for capturing nuanced social signals.
Fig.5 illustrates SocialC3D’s performance across three downstream tasks with full modality and with each of the three localized modalities ablated. The Attitude and Intent tasks show performance degradation when facial information is removed. Facial orientation offers cues about a user’s attention, while expressions convey emotional states that differentiate social behaviours with similar body movements in early stages, such as ‘throw’ and ‘wave hands’. The Action task is more sensitive to the absence of hand gestures, as detailed hand pose information helps the model to distinguish between fine-grained actions, such as the differences in finger closure for ‘touch’ and ‘handshake’. For the Intent task, the lack of gaze information leads to the greatest drop in accuracy, since gaze direction directly reflects user focus and intention. While facial orientation may offer partial substitutes, it cannot fully replicate the precision of direct gaze signals. Through knowledge distillation, the impact of missing modalities observed in the teacher model is transferred to the student model, as shown in Fig.5. These results confirm that the proposed knowledge distillation framework successfully transfers the integrated multimodal social knowledge from the teacher model to the unimodal student model.
Distillation Method JPL-Social HARPER
Intent Acc. Attitude Acc. Action Acc. Intent Acc. Attitude Acc. Action Acc.
w/o KD 80.95 59.52 59.52 61.54 28.84 42.31
Soft Label (Hinton et al., 2015) 82.14 71.43 64.29 61.54 39.13 44.23
Attention Transfer (Zagoruyko and Komodakis, 2017) 84.52 73.81 69.05 63.46 39.13 48.08
FitNet (Romero et al., 2014) 84.52 76.19 66.67 65.38 43.48 48.08
KDGAN (Wang et al., 2018) 88.10 78.57 69.05 65.38 43.48 50.00
InfoNCE (Oord et al., 2018) 88.10 80.95 75.00 67.31 52.14 57.69
Table 2. Comparison of different knowledge distillation methods when 30 % of joints and frames are corrupted.
Distillation Methods.  In Table 2, we compare the performance improvements brought by different knowledge distillation methods to the student model, SocialEgoMobile, when 30% of joints and frames are corrupted. Distillation based on intermediate features outperforms soft label distillation, as the student model in our experiments must learn to integrate information of additional modalities from high-level social representation provided by the teacher model and compensate for the impact of missing or noisy inputs. Learning only from the teacher’s output distributions is insufficient. Among the intermediate feature-based methods, InfoNCE (Oord et al., 2018) achieves the best performance. By maximizing mutual information between the teacher and student model’s intermediate social representations, InfoNCE encourages the student to learn semantically aligned features, rather than enforcing exact numerical matching at lower levels. This is suitable for our experiment setting, where the teacher and student operate on different input modalities and the student receives incomplete and noisy inputs, making low-level alignment impractical. InfoNCE promotes the alignment of high-level semantic information, leading to better noise tolerance and generalization.
Model Architecture JPL-Social
Intent F1 Attitude F1 Action F1
SocialEgoMobile 88.10 80.95 75.00
w/ 3 layers GAT 84.52 76.19 69.05
w/ 1 layers GAT 83.33 69.05 67.86
w/ 2 layer GCN (Kipf and Welling, 2017) 85.71 71.43 67.86
w/ 2 layers Bi-LSTM 88.10 82.14 75.00
w/ 1 layer TCN (Lea et al., 2017) 83.33 69.05 65.48
w/ 1 layer Transformer (Vaswani et al., 2017) 77.38 66.67 54.76
Table 3. Performance Comparison of SocialEgoMobile with Different Backbone Models on JPL-Social when 30% joints and frames are corrupted. The best results are shown in bold and the second best results are underlined.
Student Model Architecture.  To enhance the robustness and efficiency of processing noisy pose feature sequences, we designed the student model SocialEgoMobile with a two-layer GAT (Veličković et al., 2018) followed by a one-layer Bi-LSTM (Graves and Schmidhuber, 2005). Table 3 presents a comparison of different model configurations under the knowledge distillation with corrupted inputs. In the spatial dimension, a two-layer GAT achieves a good trade-off between expressiveness and stability. Fewer layers limit the receptive field, reducing the model’s ability to capture global graph structural patterns, while deeper GATs tend to propagate noise more widely. Moreover, GAT’s adaptive attention mechanism enables selective aggregation of useful information from neighbouring nodes, offering greater resilience to noisy or missing joints compared to standard GCNs (Kipf and Welling, 2017). In the temporal dimension, a single-layer Bi-LSTM provides sufficient modelling capacity for short pose feature sequences, while deeper Bi-LSTMs introduce additional parameters with marginal performance improvement. Compared to TCNs (Lea et al., 2017) and Transformers (Vaswani et al., 2017), Bi-LSTM is better suited for sequential data with inherent temporal dependencies and benefits from its gating mechanism, which improves robustness to sparsity and temporal discontinuity. Overall, the combination of two GAT layers and one Bi-LSTM in SocialEgoMobile offers an effective balance between expressiveness, robustness, and computational efficiency.
4.6. Key Takeaways
To summarize our results, we highlight three main observations:
  • The proposed teacher model SocialC3D effectively integrates multimodal inputs (body, face, hand, gaze and raw image) for egocentric social understanding, surpassing SOTAs. Each modality contributes unique and complementary information.
  • The knowledge distillation framework enables the student model to absorb multimodal and clean knowledge from the teacher, improving its ability of capture meaningful information and robustness against spatial-temporal noisy data.
  • The distilled student SocialEgoMobile achieves a trade-off between performance, robustness, and computational cost, which is suitable for deployment on resource-constrained systems.
5. Conclusion
In this paper, we address the challenges of performance, robustness, and computational efficiency in forecasting users’ social interaction intentions and behaviours from an egocentric perspective on mobile platforms or embedding system. We propose SocialEgoMobile, a lightweight social cue understanding model based on knowledge distillation. In our knowledge distillation framework, the teacher model, SocialC3D, uses multimodal inputs, including body poses, raw images, facial expressions, hand gestures, and gaze patterns, while the student model only relies on deliberate corrupted body pose features. By distilling knowledge from the teacher model’s integrated multimodal representations and downstream task supervision, the student model learns to produce robust predictions under noisy and incomplete input conditions. This work offers an effective and lightweight solution for real-time social understanding, particularly suited for resource-constrained scenarios such as social robots and intelligence social agents. Future work will focus on further enhancing the model’s robustness through real-world deployment and investigating its generalization to complex scenarios, including domain adaptation across viewpoints and environments.
This paper has been submitted to ACM Multimedia 2025.

### v2

@@
 Abstract.
-The need for social robots and agents to interact and assist humans is growing steadily.
-To be able to successfully interact with humans, they need to understand and analyse socially interactive scenes from their (robot’s) perspective.
-Works that model social situations between humans and agents are few; and even those existing ones are often too computationally intensive to be suitable for deployment in real time or on real world scenarios with limited available information.
-We propose a robust knowledge distillation framework that models social interactions through various multimodal cues, yet is robust against incomplete and noisy information during inference.
-Our teacher model is trained with multimodal input (body, face and hand gestures, gaze, raw images) that transfers knowledge to a student model that relies solely on body pose.
-Extensive experiments on two publicly available human-robot interaction datasets demonstrate that the our student model achieves an average accuracy gain of 14.75% over relevant baselines on multiple downstream social understanding task even with up to 51% of its input being corrupted.
-The student model is highly efficient: it is $<1$% in size of the teacher model in terms of parameters and uses $\sim 0.5$‰ FLOPs of that in the teacher model.
-Our code will be made public during publication.
+There is a growing need for social robots and intelligent agents that can effectively interact with and support users.
+For the interactions to be seamless, the agents need to analyse social scenes and behavioural cues from their (robot’s) perspective.
+Works that model human-agent interactions in social situations are few; and even those existing ones are computationally too intensive to be deployed in real time or perform poorly in real-world scenarios when only limited information is available.
+We propose a knowledge distillation framework that models social interactions through various multimodal cues, and yet is robust against incomplete and noisy information during inference.
+We train a teacher model with multimodal input (body, face and hand gestures, gaze, raw images) that transfers knowledge to a student model which relies solely on body pose.
+Extensive experiments on two publicly available human-robot interaction datasets demonstrate that our student model achieves an average accuracy gain of 14.75% over competitive baselines on multiple downstream social understanding tasks, even with up to 51% of its input being corrupted.
+The student model is also highly efficient - less than 11% in size of the teacher model in terms of parameters and its latency is 11.9% of the teacher model.
+Our code and related data are available at github.com/biantongfei/SocialEgoMobile.
 Human-Robot Interaction, Multimodal Learning, Knowledge Distillation, Egocentric Vision
-ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing HCI theory, concepts and models
+journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3755463ccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing HCI theory, concepts and models
 Figure 1.
-Our aim is to build a lightweight and robust model to understand human-robot social interactions.
-A novel teacher-student learning framework is built to achieve this, where the teacher model uses multimodal cues (body pose, face and hand gestures, and gaze) and raw RGB frames to learn a rich representation of the social scenario.
-Through knowledge distillation, the teacher model is able to transfers the knowledge to a lightweight student model that relies solely on body pose features.
-The student model is robust against incomplete and noisy inputs.
+We aim to build a lightweight and robust model to understand human-robot social interactions.
+A novel knowledge distillation framework is built, where the teacher model uses multimodal cues (body pose, face and hand gestures, gaze and raw images) to learn a rich social representation.
+Through distillation, the lightweight student model that relies solely on corrupted body pose features learns multimodal social knowledge and is robust against incomplete and noisy inputs.
 1. Introduction
-Recent application advances in social robotics and virtual agents include personal care, navigation, and user interaction (Zhou et al., 2022).
+Recent application advances in social robotics and agents include personal care, navigation, and user interaction (Zhou et al., 2022; Zachiotis et al., 2018; Blair et al., 2025).
 A central challenge in social AI and human-robot interaction (HRI) is enabling the systems to understand and forecast human social behaviours from an egocentric perspective, such as intent to interact, attitudes, and social actions (Bian et al., 2025).
-This capability is particularly critical during the early stages of interaction, where understanding the user’s engagement intent allows the system to proactively prepare for smooth and natural interactions.
+This capability is particularly critical during the early stages of interaction, where understanding the user’s interaction intent allows the system to proactively prepare for natural and timely interactions.
 Early prediction enhances user experience and supports adaptive, context-aware responses (Avelino et al., 2021).
 However, deploying such predictive models in real-world settings remains challenging due to constraints in computational resources, input quality, and the need for real-time processing.
-Understanding human social behaviours relies on multiple non-verbal signals (Urakami and Seaborn, 2023), including body, face and hand gestures, gaze and raw images, each providing unique and complementary information into social interactions (Yumak and Magnenat-Thalmann, 2015).
-Despite progress in understanding multimodal nonverbal cues, several challenges remain in using them for broader social understanding in real-world deployment.
-(1) Multimodal Integration: Social understanding needs fusing information from diverse modalities that can sometimes be inconsistent.
+Understanding human social behaviours relies on multiple non-verbal signals (Urakami and Seaborn, 2023), each providing unique and complementary information into social interactions (Yumak and Magnenat-Thalmann, 2015).
+Despite progress in modelling multimodal cues, several challenges remain in using them for social understanding in real-world deployment.
+(1) Multimodal Integration: Social understanding needs fusing information from diverse modalities that sometimes are inconsistent.
 Models need to integrate information and differences across modalities to extract coherent semantic representations.
-(2) Real-World Challenges: Egocentric perspectives introduce challenges such as limited field of view, camera motion, occlusions, and feature estimation errors, which degrades the reliability of social understanding.
+(2) Real-World Challenges: Egocentric perspectives introduce challenges such as limited field of view, camera motion and occlusions, which degrade the reliability of social understanding.
 (3) Computational Efficiency: Using multiple modalities requires computationally intensive models and more feature estimation modules, posing difficulties for real-time inference on resource-constrained systems, such as social robots.
-Addressing these challenges needs a lightweight yet robust framework that can integrate multimodal cues under noisy conditions while supporting real-time social understanding.
-To tackle these challenges, we propose a knowledge distillation framework that enables a lightweight student model to infer rich social cues from incomplete body pose input alone and jointly forecast three key aspects of social understanding: interaction intentions, attitudes, and social actions, as shown in Fig.1.
-Our approach consists of two stages: first, we train a teacher model with full-modal inputs, including body, face and hand gestures, gaze and raw images, to learn and output high-quality social representations.
-Then, we distil this knowledge into a lightweight student model that relies solely on body pose while incorporating spatio-temporal noise to simulate real-world data degradation.
-Instead of only learning from downstream task labels, the student also learns to infer multimodal and clean social cues through knowledge distillation, ensuring robust performance under noisy conditions.
-The lightweight design of the student model improves computational efficiency.
-In summary, our work enhances the robustness and efficiency of jointly forecasting users’ intent to interact, attitudes, and social actions from an egocentric perspective through the contributions:
+To tackle these challenges, we propose a knowledge distillation framework that enables a lightweight student model to infer rich social cues from corrupted body pose input and jointly forecast three key aspects of social understanding: intent to interact, attitudes, and social actions, as shown in Fig.1.
+Our approach consists of two stages: first, we train a teacher model with multimodal inputs, including body, face and hand gestures, gaze and raw images, to learn high-quality social representations.
+Then, we distil this knowledge into a lightweight student model that relies solely on body pose while containing spatiotemporal noise to simulate real-world data degradation, ensuring robust performance under noisy conditions.
+Body pose features are crucial cues for social understanding (Bian et al., 2025) and rely only on lightweight feature extraction and analysis models.
+In summary, our work enhances the robustness and efficiency of egocentric social understanding through the contributions:
   • 1
@@
   • 2
-    Robustness to real-world noises: To enhance robustness against noisy and incomplete input, we introduce spatiotemporal corruption for the student model.
-With knowledge distillation, when 51% of spatiotemporal information is lost the student maintains an average accuracy of 81.35% on JPL-Social (Bian et al., 2025) (14.68% improvement over independent training) and 59.05% on HARPER (Avogaro et al., 2024) (14.82% improvement).
+    Robustness to noises: To enhance robustness against noisy and incomplete input, we introduce spatiotemporal corruption for the student model during distillation.
+When 51% of spatiotemporal information is lost, the student maintains an average accuracy of 81.35% on JPL-Social (Bian et al., 2025) (14.68% improvement over independent training) and 59.05% on HARPER (Avogaro et al., 2024) (14.82% improvement).
   • 3
-    Efficiency and deployment readiness: The proposed student model SocialEgoMobile demonstrates high computational efficiency, requiring only 1% of model parameters and 0̃.5‰ of the floating-point operations (FLOPs) compared to the teacher model.
+    Efficiency and deployment readiness: The proposed student model SocialEgoMobile demonstrates high computational efficiency, requiring only 1% of model parameters and 11.9% latency compared to the teacher model SocialC3D.
 2. Related Work
 2.1. Social Intent Prediction
-Egocentric social intent prediction is crucial for socially intelligent agents to establish natural, harmonious, and efficient interactions with users.
-A precise understanding of the user’s social behaviour enables agents to timely adjust response strategies to meet user needs.
+Egocentric social intent prediction is crucial for social robots to establish natural, harmonious, and efficient interactions with users.
+A precise understanding of the user’s social behaviour enables robots to timely adjust response strategies to meet user needs.
 Current research primarily models social intent as a binary classification task, focusing on predicting whether a user intends to interact with an intelligent agent.
@@
 2.2. Non-Verbal Cues in Multimodal HRI
-In Human-Robot Interaction (HRI), accurately interpreting the user’s identity, social intent, behaviours, and communication requires integrating information from multiple modalities (Yumak and Magnenat-Thalmann, 2015).
-Meanwhile, non-verbal cues play a crucial role in understanding user intentions and needs (Urakami and Seaborn, 2023).
-For instance, while detecting sarcasm through text is challenging during a conversation, recognizing subtle body movements such as shoulder shrugs or lip curls allows humans to accurately identify sarcasm (Castro et al., 2019).
+In Human-Robot Interaction (HRI), accurately interpreting the user’s identity, social intent and behaviours requires integrating multimodal cues (Yumak and Magnenat-Thalmann, 2015).
+Meanwhile, non-verbal cues play a crucial role in social understanding (Urakami and Seaborn, 2023).
+For instance, while detecting sarcasm through text is challenging, recognizing subtle body movements such as shoulder shrugs allows humans to accurately identify sarcasm (Castro et al., 2019).
 Consequently, capturing and interpreting mutlimodal non-verbal cues in HRI is essential for assessing the users.
@@
 Research gap.
-Despite the progress in social understanding using deep learning models, existing approaches often rely on multimodal inputs.
-However, processing more modal information increases computational cost (Liang et al., 2024) and the deep models are sensitive to missing or noisy information (Drenkow et al., 2021), making them impractical for deployment on resource-constrained devices and real-world scenarios.
-While knowledge distillation has been explored to compress large models into lighter versions, most prior works focus on preserving accuracy under ideal conditions, overlooking the need to integrate multimodal information using a single modality and maintain robustness in environments with incomplete or degraded inputs (Arani et al., 2021).
-To address these limitations, we propose a knowledge distillation framework with a lightweight, body pose-only student model that learns from a full-modality teacher model through a robust distillation, ensuring both computational efficiency and resilience to temporal and spatial noise.
+While deep learning models using multimodal cues have advanced social understanding, their high computational demands (Liang et al., 2024) and sensitivity to noisy or incomplete inputs (Drenkow et al., 2021) limit deployment on resource-constrained platforms.
+Existing knowledge distillation approaches primarily aim to retain accuracy under ideal conditions (Arani et al., 2021), neglecting robustness and integrating multimodal information using a single modality.
+To overcome these issues, we propose a robust distillation framework where a lightweight, body pose-only student model learns from a multimodal teacher, achieving efficient, resilient, and accurate social understanding.
 3. Methodology
 Figure 2.
-Our knowledge distillation framework uses SocialC3D as the teacher model, which fuses raw images, body, face, hand gestures, and gaze.
-Each modality is processed by a ResNet (He et al., 2016) and integrated via lateral connections and late fusion, producing a high-quality social representation for downstream tasks.
-The lightweight student model, SocialEgoMobile, uses only corrupted body pose.
-It consists of a two-layer GAT (Veličković et al., 2018) and a single-layer Bi-LSTM (Graves and Schmidhuber, 2005) to extract social representations.
-The framework distillates knowledge from the teacher model by maximising the mutual information of the social representations output by the teacher and student model.
+Our knowledge distillation framework uses SocialC3D as the teacher model, which fuses raw images, body, face, hand gestures, and gaze information, producing a multimodal social representation.
+The lightweight student model, SocialEgoMobile, uses only body pose features to output social representations.
+The framework maximises the mutual information of the social representations of the teacher and student model, to transfer social knowledge.
 3.1. Problem Definition
-Given a sequence of egocentric videos captured by an intelligent social agent, our aim is to forecast the user’s social intent and behaviours in real time based only on the early stages of interaction (e.g., the first second).
-Specifically, we forecast the user’s intent to interact, attitude, and social actions.
-To achieve this, we propose a knowledge distillation framework where a lightweight student model is trained to rely on body pose features, while a teacher model uses full multimodal inputs (body, face, hand gesture, gaze and raw images), as shown in Fig.2.
+Given a sequence of egocentric videos captured by a social robot, we aim to forecast the user’s social intent, attitude and action in real-time by observing the early stage (the first second).
+So we propose a knowledge distillation framework where a lightweight student model is trained to rely on body pose features, while a teacher model uses multimodal inputs (body, face, hand gesture, gaze and raw images), as shown in Fig.2.
 Body pose features, facial keypoints, and hand keypoints are extracted from the video using Alphapose (Fang et al., 2022) and gaze features are extracted using MCGaze (Guan et al., 2023).
@@
 The teacher model $F_{teacher}$ is trained with the complete set of input modalities:
-(1) $R_{teacher}=F_{teacher}(\boldsymbol{x}_{t:T}^{\text{Pose}},\boldsymbol{x}_{t:T}^{\text{Image}},\boldsymbol{x}_{t:T}^{\text{Face}},\boldsymbol{x}_{t:T}^{\text{Hand}},\boldsymbol{x}_{t:T}^{\text{Gaze}})$
-where $R_{t}$ represents the high-level social representation extracted by the teacher model.
-To ensure efficiency and robustness, we introduce a lightweight student model $F_{student}$, which relies solely on body pose features $\boldsymbol{x}_{t}^{\text{Pose}}$:
-(2) $R_{student}=F_{student}(\boldsymbol{x}_{t:T}^{\text{Pose}})$
-where $R_{student}$ represents the social representation extracted by the student model.
-For the downstream tasks of forecasting user’s intent to interact, attitude, and social actions, the social representations $R$ extracted by the teacher or student model are processed by the Chain hierarchical classifier (Bian et al., 2025), which simulates how human understand and forecast social signals, predicting the results of three downstrem tasks jointly:
+(1) $R_{teacher}=F_{teacher}(\boldsymbol{x}_{t}^{\text{Pose}},\boldsymbol{x}_{t}^{\text{Image}},\boldsymbol{x}_{t}^{\text{Face}},\boldsymbol{x}_{t}^{\text{Hand}},\boldsymbol{x}_{t}^{\text{Gaze}}),t\in[0,T]$
+where $R_{teacher}$ represents the high-level social representation outputed by the teacher model and $T$ represents the length of the input sequence.
+The lightweight student model $F_{student}$ relies solely on body pose features $\boldsymbol{x}_{t}^{\text{Pose}}$:
+(2) $R_{student}=F_{student}(\boldsymbol{x}_{t}^{\text{Pose}}),t\in[0,T]$
+where $R_{student}$ represents the social representation outputed by the student model.
+For the subtasks of forecasting user’s intent to interact, attitude, and social actions, the social representations $R$ extracted by the teacher or student model are processed by a Chain hierarchical classifier (Bian et al., 2025), which simulates how human infer others’ social intents, predicting the results of three downstream tasks jointly:
 (3) $\mathcal{L}_{intent},\mathcal{L}_{attitude},\mathcal{L}_{action}=Classifier(R)$
 3.2. Teacher Model
-To effectively integrate multiple modalities, we employ a teacher model, SocialC3D which is based on PoseC3D (Duan et al., 2022), a spatio-temporal feature extraction framework using 3D convolutional neural networks (3D-CNN).
+To integrate multiple modalities, we employ a teacher model, SocialC3D which is based on PoseC3D (Duan et al., 2022), a spatio-temporal feature extraction framework using 3D convolutional neural networks (3D-CNN).
 PoseC3D converts sequences of 2D pose keypoints into 3D heatmap stacks, according to their coordinates and confidence scores.
@@
 (4) $H_{i}(x,y)=\exp\left(-\frac{(x-x_{i})^{2}+(y-y_{i})^{2}}{2\sigma^{2}}\right)*c_{i}$
-where $\sigma$ is the Gaussian variance of the heat map; ($x$, $y$) are the coordinates of the heatmap; ($x_{i}$, $y_{i}$) and $c_{i}$ are the coordinates and confidence scores of the current keypoint $i$.
-This approach effectively captures spatio-temporal dependencies and exhibits greater robustness to noise than traditional skeleton-based methods.
-Additionally, the lateral connections between the RGB and Pose pathway fuse image and pose information, enhancing multimodal learning.
-Based on this, we introduce SocialC3D to incorporate more modal information.
-SocialC3D transforms additional modalities, facial expressions, hand gestures, and gaze, into 3D heatmap stacks as input, which are processed using the same ResNet (He et al., 2016) as the body pose stream.
-For gaze estimation, we use the centre of the head bounding box as the beginning and calculate the end based on the gaze direction and the diagonal length of head bounding box.
+where $\sigma$ is the Gaussian standard deviation of the heatmap; ($x$, $y$) are the position in the heatmap; ($x_{i}$, $y_{i}$) and $c_{i}$ are the coordinates and confidence scores of the keypoint $i$.
+This approach exhibits greater robustness to pose estimation error than skeleton-based methods (Duan et al., 2022).
+Also, the lateral connections between the RGB and Pose pathway fuse information on multiple levels.
+Based on this, we introduce SocialC3D to incorporate additional modalities, including facial expressions, hand gestures, and gaze, into 3D heatmap stacks as input, which are processed using the same ResNet structure (He et al., 2016) as the body pose pathway.
+For gaze vector, we use the centre of the head bounding box as the beginning and calculate the end based on the gaze direction and the diagonal length of head bounding box.
 SocialC3D incorporate lateral connections to propagate intermediate representations across modalities at multiple stages, enabling deeper feature integration.
+After the ResNets, the outputs of different modalities are concatenated and passed through a linear layer to produce a 16-dimensional multimodal social representation.
 3.3. Student Model
-The student model takes only body pose keypoints as input.
-Compared to raw visual information, body language provides a more direct and interpretable signal for understanding social interactions.
-Additionally, body posture inherently contain some facial and hand dynamics, allowing the model to integrate multimodal cues through knowledge distillation.
-Moreover, body pose features, as high-level abstractions, support lighter spatio-temporal analysis models and pose estimation, enhancing the student model’s suitability for resource-constrained settings.
-To reduce computational complexity while maintaining strong predictive performance, we developed SocialEgoMobile as our student model.
-To reduce the risk of overly deep networks amplifying noise and hindering useful information extraction, we designed a streamlined architecture consisting of a two-layer graphical attention network (GAT) (Veličković et al., 2018) and a single-layer bi-directional long- and short-term memory (Bi-LSTM) (Graves and Schmidhuber, 2005).
-By dynamically adjusting the attentional weights between neighbouring joints, GAT can effectively capture the spatial relationships between key points of the body pose skeleton graph, thus making the model more focused on useful information and limiting the noise.
-Meanwhile, Bi-LSTM enhances temporal modelling by processing sequences forward and backward with forgetting gates that can selectively filter noise.
-This lightweight yet effective design maintains stable performance under noisy and incomplete input conditions, thus enhancing robustness and making it suitable for real-time deployment in resource-limited environments.
+The student model takes only body pose features as input.
+Compared to raw visual information, body language provides a more direct and interpretable cue for social understanding and body pose is one of the most effective modalities.
+Additionally, body posture inherently contains some facial and hand dynamics, allowing the model to integrate multimodal cues through knowledge distillation.
+Moreover, body pose, as a high-level abstraction, supports lighter analysis models and pose estimation methods, enhancing the student model’s suitability for resource-constrained settings.
+To achieve a balance between performance, robustness to noise and computational efficiency, we developed SocialEgoMobile as our student model, consisting of a two-layer graphical attention network (GAT) (Veličković et al., 2018) and a single-layer bi-directional Long Short Term Memory (Bi-LSTM) (Graves and Schmidhuber, 2005).
+By dynamically adjusting the attentional weights between neighbouring joints, GAT can capture the spatial relationships of the body pose skeleton for each frame, focusing on useful information and limiting the noise.
+Meanwhile, Bi-LSTM captures the temporal dynamic of the input sequence forward and backward and filters noise using forgetting gates.
+The hidden dimension of the GATs is 16, and that of the Bi-LSTM is 128.
+The outputs are projected into a 16-dimensional social representation via a linear layer.
+This lightweight yet effective design maintains stable performance under noisy input conditions, thus enhancing robustness and making it suitable for deployment in resource-limited environments.
 3.4. Knowledge Distillation
-To effectively transfer knowledge of the user’s social behaviour from the teacher model to the student model, we adopt a feature-based knowledge distillation method.
-The student model learns not only from the ground truth labels of the downstream tasks but also by minimizing the discrepancy between its high-level social representations and those of the teacher model.
-We employ InfoNCE (Information Noise Contrastive Estimation) (Oord et al., 2018) to estimate and maximize the mutual information between teachers’ and students’ social representations.
-Specifically, InfoNCE enhances similarity between social representation of teacher and student models of the same input while reducing similarity between different inputs.
-During training, the loss function of the student model combines the mutual information loss from InfoNCE with the classification loss from the downstream tasks, weighted for back propagation.
-The loss of InforNCE is defined as:
+To transfer social knowledge from the teacher model to the student model, we adopt a feature-based distillation method, InfoNCE (Information Noise Contrastive Estimation) (Oord et al., 2018).
+The student learns from the ground truth labels of the downstream tasks and the difference between its social representation and that of the teacher.
+InfoNCE estimates and maximises its mutual information by enhancing the similarity between the social representations of teacher and student models of the same video clip while reducing the similarity between different ones.
+The distillation loss is defined as:
 (5) ${{Loss}_{\text{infoNCE}}}=-\mathop{{}\mathbb{E}}\left[\log\frac{\exp(f(x,x^{+}))}{\sum_{x_{j}\in X}\exp(f(x,x_{j}))}\right]$
-where $x$ is the teacher model’s social representation for a given input, $x^{+}$ is the representation from the student model for the same input, and $x_{j}$ is the representation from the student model for different inputs.
+where $x$ is the teacher’s social representation for a given video clip, $x^{+}$ is the representation from the studen for the same clip, and $x_{j}$ is the representation from the student for other clips.
 The similarity function $f(x,x^{\prime})$ is defined using cosine similarity as follows:
 (6) $\mathbf{}{f}(\mathbf{x},\mathbf{x^{\prime}})=\frac{\mathbf{x}\cdot\mathbf{x^{\prime}}}{\|\mathbf{x}\|\|\mathbf{x^{\prime}}\|}$
-By knowledge distillation, the student model can still learn high-level social knowledge from the teacher model with limited modality and corrupted inputs.
-Furthermore, we introduce spatio-temporal corruption in the student model’s input, enhancing its ability to integrate multimodal social cues while maintaining robustness to noise and missing data.
 3.5. Input Corruption
-In real-world applications, the social understanding model must not only work under the computational constraints of mobile platforms but also remain robust to incomplete and noisy inputs.
-To simulate such conditions, we deliberately introduce corruption in both spatial and temporal dimensions during training.
-In the spatial domain, a fixed number of body pose keypoints are randomly selected and corrupted by either setting their coordinate values and confidence scores to zero or replacing the coordinates with random values within the normalized range.
-This mimics localized body occlusions or pose estimation errors.
-In the temporal domain, we randomly drop a fixed number of frames by setting all keypoint values in those frames to zero or normalized random noise, simulating transient global occlusions.
-These corruptions force the student model to learn robust representations and maintain performance under incomplete or noisy input during knowledge distillation.
+In real-world applications, social understanding models must work under the computational constraints of mobile platforms and be robust to incomplete and noisy inputs.
+To simulate such conditions, we deliberately introduce corruption in spatial and temporal dimensions.
+In the spatial domain, a fixed number of body pose keypoints are randomly corrupted by setting their coordinate values and confidence scores to zero or replacing the coordinates with normalized random values.
+This mimics localized body occlusions and pose estimation errors.
+In the temporal domain, we randomly corrupt frames by setting all keypoint values to zero or applying normalised random noise, simulating transient full-body occlusions.
+These corruptions force the student model to learn robust representations and maintain performance under incomplete and noisy input.
 4. Experiments
 4.1. Data
-Figure 3. We standardized the input frame rate of two datasets to 10 FPS and limited the observation window to the first second of social interaction.
-To evaluate the performance and robustness of our models in forecasting users’ social intents and behaviours from an egocentric perspective, we use two datasets of human-robot social interactions recorded from the robot’s perspective.
-Examples of datasets are shown in Fig.
+Figure 3. We set the FPS of two datasets to 10 and limited the observation window to the first second of social interactions.
+To evaluate the performance and robustness of our models in egocentric social understanding, we use two datasets of human-robot social interactions recorded from the robot’s perspective.
+Examples of the datasets are shown in Fig.
 3.
-JPL-Social (Bian et al., 2025) is an extension of the JPL-Interaction datasets (Ryoo and Matthies, 2013; Ryoo et al., 2015), designed for human-robot interaction research, with a focus on users’ social actions within the central visual field.
-The original dataset captures 8 social actions performed by the main user: handshake, hug, pet, wave, punch, throw, point, and leave.
-JPL-Social extends this by annotating additional users in the field of view during multi-person scenarios, introducing two additional social actions: gaze and no response.
-Furthermore, JPL-Social provides labels for social intent (categorized as interested, interacting, or not interested) and attitude (positive or negative) for each user.
+JPL-Social (Bian et al., 2025) is an extension of the JPL-Interaction datasets (Ryoo and Matthies, 2013; Ryoo et al., 2015), focused on the user’s social actions.
+The original dataset captures 8 social actions performed by the users in the centre of the screen: handshake, hug, pet, wave, punch, throw, point, and leave.
+JPL-Social extends this by annotating additional users in the background during multi-person scenarios, introducing two additional action label: gaze and no response.
+Furthermore, JPL-Social provides annotations for social intent (categorized as interacting, interested and not interested) and attitude (positive or negative) for each user.
 The dataset comprises 290 interaction sequences involving 16 participants in diverse environmental settings.
-HARPER (Avogaro et al., 2024) is a dataset capturing 10 human-robot social actions performed by 17 users, recorded from multiple perspectives using five cameras mounted on a quadrupedal robot.
-For our research, we use 6 social actions specifically relevant to robot-oriented interactions: crash, walk-avoid, touch, walk-stop, punch, and kick, so there are 260 interaction sequences.
-In addition, HARPER includes variations in viewing angle and resolution, increasing the difficulty of understanding users’ social signals.
-To ensure consistency in social representation, we manually annotated HARPER with the same social intent and attitude labels as JPL-Social.
+HARPER (Avogaro et al., 2024) is a dataset recorded from multiple perspectives using five cameras mounted on a quadrupedal robot.
+We use 6 social actions relevant to robot-oriented interactions: crash, walk-avoid, touch, walk-stop, punch and kick, so there are 260 sequences.
+In addition, HARPER includes variations in viewing angle and resolution, increasing the difficulty of understanding users’ behaviours.
+Following the protocol of JPL-Social (Bian et al., 2025), we manually annotate HARPER for the intent and attitude prediction tasks.
 4.2. Experiment Setting
 We set the Frames Per Second (FPS) of the videos to 10 and limited the observation window to the first second of social interaction, which starts when the user approaches the social agent, so the inputs have 10 frames.
-To expand the training set, we randomly cropped video frames and re-extracted pose features, facial expressions, hand gestures, and gaze information.
-After horizontal flipping, we augmented number of videos 20 times for richer training data.
+To expand the training set, we randomly cropped and horizontally flipped video frames, augmenting the number of videos 20 times for richer training data.
 We train and test the models on the two datasets separately.
-Given our focus on model efficiency, we report both the number of model parameters and the number of floating point operations (FLOPs) for each input.
+Given our focus on model efficiency, we report the number of model parameters and the average latency for each input.
+All experiments were conducted on an NVIDIA RTX 4070 GPU.
+All models are trained for 40 epochs with a batch size of 96.
+Training is early stopped if the average performance across the three subtasks does not improve for 5 epochs.
+Each modality channel in the SocialC3D is trained independently first and then jointly fine-tuning.
+To enhance performance, the body pose and raw image channels of all teacher models are initialized with pre-trained weights on Kinetics 400 (Kay et al., 2017).
+SocialC3D is optimized using SGD (Sutskever et al., 2013), while SocialEgoMobile uses the Adam optimizer (Kingma and Ba, 2015).
+During knowledge distillation, the total loss is computed as a weighted sum of the classification loss and the distillation loss, with weights of 0.8 and 0.2.
+The temperature for the InfoNCE (Oord et al., 2018) is set to 0.1.
 4.3. Performance of the Teacher Model
-Model Params FLOPs JPL-Social HARPER
-(M) (M) Intent Acc. Attitude Acc. Action Acc. Intent Acc. Attitude Acc. Action Acc.
-ST-GCN+ (Yan et al., 2018) 43.86 1924.87 86.90 76.19 71.43 86.54 73.08 78.85
-ST-TR+ (Plizzari et al., 2021) 58.48 4724.63 79.76 59.52 48.81 75.00 65.38 59.61
-MS-G3D+ (Liu et al., 2020) 48.82 5241.97 88.10 80.95 76.19 90.38 78.85 80.77
-SocialEgoNet+ (Bian et al., 2025) 37.78 422.35 86.90 77.38 71.43 86.54 75.00 78.85
-SocialC3D (ours) 48.49 7902.19 92.85 88.10 82.14 96.15 82.69 88.46
-SocialEgoMobile (ours) 0.43 4.23 82.14 71.43 67.86 69.23 44.23 51.92
 Table 1.
-Comparison of SocialC3D and SocialEgoMobile with state-of-the-art methods on the JPL-Social and HARPER.
-SocialEgoMobile relies solely on clean body pose features as input.
-’+’ indicates that the model uses raw image and gaze information.
+Comparison of SocialC3D and SocialEgoMobile with state-of-the-art.
+’+’ indicates that the model uses additional modalities same as SocialC3D.
+$\Delta_{1}$ denotes the time to extract whole-body pose and gaze features from a single frame, which is 4.96 ms under our experimental setup.
+$\Delta_{2}$ refers to the extraction time for body pose features, which is 3.06 ms.
+Model Params Latency JPL-Social HARPER
+Intent Attitude Action Intent Attitude Action
+(M) $\downarrow$ (ms) $\downarrow$ Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
+ST-GCN+ (Yan et al., 2018) 43.86 $\Delta_{1}$ + 10.31 86.90 85.93 76.19 74.56 71.43 65.23 86.54 84.02 73.08 74.68 78.85 77.18
+ST-TR+ (Plizzari et al., 2021) 58.48 $\Delta_{1}$ + 14.33 79.76 77.82 59.52 58.87 48.81 46.76 75.00 66.37 65.38 58.91 59.61 39.63
+MS-G3D+ (Liu et al., 2020) 48.82 $\Delta_{1}$ + 13.74 88.10 88.11 80.95 78.12 76.19 74.59 90.38 88.11 78.85 76.52 80.77 74.00
+SocialEgoNet+ (Bian et al., 2025) 37.78 $\Delta_{1}$ + 10.05 86.90 86.54 77.38 75.68 71.43 68.93 86.54 77.29 75.00 73.72 78.85 71.78
+SocialC3D 48.49 $\Delta_{1}$ + 22.34 92.85 91.25 88.10 88.91 82.14 78.09 96.15 96.14 82.69 88.28 88.46 88.35
+SocialEgoMobile 0.43 $\Delta_{2}$ + 0.19 86.90 85.26 71.43 68.61 67.86 62.63 69.23 66.92 38.46 42.07 51.92 46.56
 As the teacher model, we compare SocialC3D with several state-of-the-art methods.
-Although originally designed for body pose skeleton graphs, these methods can also process facial and hand keypoint graphs.
-To ensure a fair comparison, we incorporated the raw image and gaze information channels from SocialC3D into all teacher models, aligning the available modalities across teacher models.
-Each modality channel in the teacher models is trained independently first and then during joint fine-tuning, outputs of all the channels are fused before being passed to the chain hierarchical classifier.
-To enhance body language understanding, the body pose and raw image channels of all teacher models are initialized with pre-trained weights on Kinetics 400 (Kay et al., 2017).
-As shown in Tab le1, we compare SocialC3D with SOTA methods on the JPL-Social and HARPER datasets.
-SocialC3D consistently achieves superior performance on three downstream social tasks in both datasets.
-Compared with the second-best model, MS-G3D (Liu et al., 2020), SocialC3D achieves an average accuracy improvement of 5.85% on JPL-Social and 5.77% on HARPER, which highlights the effectiveness of SocialC3D in understanding social cues and forecasting users’ social intentions and behaviours from an egocentric perspective.
+Although originally designed for pose skeletons, these methods can also process facial and hand keypoints.
+To ensure a fair comparison, we incorporated the raw image and gaze information channels from SocialC3D into all teacher models, aligning the modalities across teacher models.
+In Table 1, SocialC3D consistently achieves superior performance on three downstream social tasks on both datasets.
+Compared with the second-best model, MS-G3D (Liu et al., 2020), SocialC3D achieves an average accuracy improvement of 5.85% on JPL-Social and 5.77% on HARPER, which highlights the effectiveness of SocialC3D in understanding social cues and forecasting users’ social intentions from an egocentric perspective.
 We attribute this improvement to two key factors.
 First, instead of directly using pose features, converting them into heatmaps enhances robustness to pose estimation errors, which is particularly pronounced in egocentric settings due to narrow and jittery fields of view.
-Second, our experiments focus on the initial 10 frames of social interactions.
-3D CNNs are suitable for capturing local temporal dynamics, making them more appropriate for our experiment setting.
-In contrast, models like ST-GCN (Yan et al., 2018) and MS-G3D (Liu et al., 2020) rely on complete and accurate pose features, making them more vulnerable to missing joints and estimation noise.
-ST-TR (Plizzari et al., 2021), while effective in modelling long-term dependencies due to using transformers (Vaswani et al., 2017), struggles with short sequences and lacks mechanisms to handle input degradation.
-Although SocialEgoNet (Bian et al., 2025) is designed to capture users’ non-verbal social cues from an egocentric perspective, its lightweight architecture limits its capacity to extract fine-grained information independently.
+Second, our experiments focus on the first 10 frames of social interactions.
+3D CNNs are suitable for capturing short-term temporal dynamics, making them more appropriate for our experiment setting.
+In contrast, models like ST-GCN (Yan et al., 2018) and MS-G3D (Liu et al., 2020) rely on complete and accurate pose features, making them more vulnerable to missing joints and estimation errors.
+ST-TR (Plizzari et al., 2021), while effective in modelling long-term dependencies due to using transformers (Vaswani et al., 2017), struggles with short sequences.
+Although SocialEgoNet (Bian et al., 2025) is designed to capture users’ non-verbal social cues from an egocentric perspective, its lightweight architecture limits its capacity to extract fine-grained information.
 In addition, we report the performance of SocialEgoMobile using only body pose to compare its accuracy and computational efficiency against the teacher models.
-Compared to SocialC3D, SocialEgoMobile’s average accuracy of downstream tasks on JPL-Social was 13.
-89% lower and 33.97% lower on HARPER.
-The larger performance gap on HARPER is attributed to its recording setup, where the robot dog has a lower vantage point and fewer frames with full-body visibility, resulting in noisier and more incomplete pose inputs, as shown in Fig.3.
-In terms of computational efficiency, compared to SocialC3D, SocialEgoMobile only needs 0.98% of the model parameters and 0.54‰of FLOPs.
-These results show SocialEgoMobile is computationally lightweight enough for social understanding on mobile platforms with limited resources.
+Compared to SocialC3D, SocialEgoMobile’s average accuracy of downstream tasks was 13.89% lower on JPL-Social and 33.97% lower on HARPER.
+The larger performance gap on HARPER is attributed to its recording setup, where the robot dog has a lower viewing angle and fewer frames with full-body visibility, resulting in noisier and more incomplete pose inputs, as shown in Fig.3.
+In such conditions, high-level features such as body pose become less reliable.
+These features rely on per-frame feature estimation methods, which lack temporal context and will produce unstable outputs, leading to fragmented feature sequences and degraded model understanding.
+In contrast, raw image sequences provide a temporally coherent context.
+Thus, teacher models with image inputs can maintain high performance on HARPER, while SocialEgoMobile, relying solely on body pose, suffers from significant degradation.
+In terms of computational efficiency, we report the number of parameters and the inference latency for all methods.
+Since the pose and gaze estimation algorithms can run at 10 FPS, the reported latency only includes the model inference time and the feature extraction time for the last frame.
+SocialEgoMobile only needs 0.43 M model parameters (0.98% of SocialC3D) and 3.25 ms latency (11.9% of SocialC3D).
+These results show that SocialEgoMobile is lightweight enough for real-time social understanding on mobile platforms with limited resources.
 4.4. Robustness Analysis on the Student Model
 Figure 4.
-Knowledge distillation (KD) consistently improves the performance of the student model, SocialEgoMobile, under Individual and simultaneous spatio-temporal corruption on all three downstream tasks, interaction intent, attitude, and social action forecast.
+Knowledge distillation (KD) consistently improves the performance of the student model, SocialEgoMobile, under spatio-temporal corruption on all three downstream tasks, interaction intent, attitude, and social action forecast.
 Improvements on downstream task accuracy through distillation are labelled.
-To address the practical challenges faced by mobile socially intelligent agents from an egocentric perspective, such as occlusion of user body parts and pose estimation error, we deliberately introduced random spatio-temporal corruption into the input of the student model SocialEgoMobile, as mention in Section 3.5.
-The goal is to improve robustness by using the rich and clean knowledge extracted from the intermediate social representations of the teacher model SocialC3D, to supervise the learning of the student model with corrupted inputs.
-During the robustness tests, student models were trained for 70 epochs.
-The first 30 epochs are warm-up phase, where the corruption rate was gradually increased to the final value to prevent unstable training due to excessive information loss.
-Fig.4 compares SocialEgoMobile’s performance under two settings: independent learning and knowledge distillation.
+To address the practical challenges faced by mobile socially intelligent agents from an egocentric perspective, such as occlusion of user body parts and pose estimation error, we deliberately introduced random spatiotemporal corruption into the input of the student model SocialEgoMobile, as mentioned in Section 3.5.
+The goal is to improve robustness by using the multimodal and uncorrupted knowledge extracted from the teacher model, SocialC3D, to supervise the learning of the student model with corrupted inputs.
+During the robustness experiments, student models were trained for 70 epochs.
+The first 30 epochs are the warm-up phase, where the corruption rate is gradually increased to the target value to prevent unstable training due to excessive information loss.
+Figure 5. Comparison of SocialC3D using different input modalities and the its impact on SocialEgoMobile (SEM). Accuracy drops due to missing modalities are labelled.
+Fig.4 compares SocialEgoMobile’s performance under independent learning and knowledge distillation.
 Under independent learning, performance degrades as corruption increases.
-When 30% of joints and frames is corrupted, $30\%\oplus 30\%=51\%$ of the information in the input was corrupted.
-The average accuracy across three downstream tasks drops by 8.85%, demonstrating the increased difficulty and robustness demands.
+When 30% of joints and frames are corrupted, $30\%\oplus 30\%=51\%$ of the information in the input was corrupted.
+This ratio maintains a balance of noise and useful information.
+The average accuracy across three subtasks drops by 8.85%, demonstrating the increased difficulty and robustness demands.
 In contrast, with knowledge distillation, SocialEgoMobile consistently outperforms the independently trained counterpart across all input settings.
-When 30% of joints and frames is corrupted, SocialEgoMobile trained via distillation surpasses the performance of the independently trained model using clean input.
-This suggests that knowledge distillation effectively transfers social knowledge from the teacher model, which incorporates additional modalities and complete, clean inputs, to the student model.
+When 30% of joints and 30% of frames are corrupted, SocialEgoMobile trained via distillation surpasses the performance of the independently trained one without input corruption.
+This suggests that knowledge distillation effectively transfers social knowledge from the teacher model, which incorporates additional modalities and complete inputs to the student model.
 So it can more accurately capture valuable social cues even when facing corrupted inputs.
@@
 This is because temporal patterns in human behaviour tend to be more regular and predictable, allowing models to more effectively compensate for missing frames and filter out noise.
-In contrast, spatial patterns of pose graphs are more complex, such as body structure and human kinematics.
+In contrast, spatial patterns of pose skeletons are more complex, such as body structure and human kinematics.
 4.5. Ablation Experiments
-Figure 5.
-Performance comparison on JPL-Social and HARPER of SocialC3D using different input modalities, and the impact of corresponding SocialC3D-based Knowledge Distillation on SocialEgoMobile (SEM) when 30% of spatio-temporal information are corrupted.
-Accuracy drops due to missing modalities are labelled.
 Contribution of Mutli-Modalities.
-We incorporate body pose, raw images, face, hand gestures, and gaze as input modalities for the teacher model SocialC3D, enabling it to comprehensively capture social cues from users’ non-verbal behaviours across multiple perspectives.
+We incorporate body pose, raw images, face, hand gestures, and gaze as input modalities for the teacher model SocialC3D, enabling it to comprehensively capture uers’ non-verbal social cues.
 Body pose and raw images are the foundational modalities and prior work has demonstrated their effectiveness and complementarity in understanding body movements (Duan et al., 2022).
 In contrast, facial expressions, hand gestures, and gaze focus on localized regions of the body, providing fine-grained cues essential for capturing nuanced social signals.
-Fig.5 illustrates SocialC3D’s performance across three downstream tasks with full modality and with each of the three localized modalities ablated.
-The Attitude and Intent tasks show performance degradation when facial information is removed.
+Fig.5 illustrates the performance of SocialC3D and SocialEgoMobile under its supervision with full modality and with each of the three localized modalities ablated.
+Removing any modality degrades the overall performance of SocialC3D, but the extent of impact varies across social sub-tasks.
+Specifically, the Attitude and Intent tasks show performance degradation when facial information is removed.
 Facial orientation offers cues about a user’s attention, while expressions convey emotional states that differentiate social behaviours with similar body movements in early stages, such as ‘throw’ and ‘wave hands’.
@@
 While facial orientation may offer partial substitutes, it cannot fully replicate the precision of direct gaze signals.
-Through knowledge distillation, the impact of missing modalities observed in the teacher model is transferred to the student model, as shown in Fig.5.
+For the student model, 30% of the frames and joints are corrupted in the input.
+We maintain this corruption ratio to ensure that the findings remain applicable to real-world deployment scenarios.
+We consider this corruption setting representative, as it strikes a balance between noise and meaningful information.
+Through knowledge distillation, the impact of missing modalities on the teacher is transferred to the student, as shown in Fig.5.
 These results confirm that the proposed knowledge distillation framework successfully transfers the integrated multimodal social knowledge from the teacher model to the unimodal student model.
+Table 2. Comparison of different knowledge distillation methods when 30 % of joints and frames are corrupted.
 Distillation Method JPL-Social HARPER
@@
 InfoNCE (Oord et al., 2018) 88.10 80.95 75.00 67.31 52.14 57.69
-Table 2. Comparison of different knowledge distillation methods when 30 % of joints and frames are corrupted.
 Distillation Methods.
-In Table 2, we compare the performance improvements brought by different knowledge distillation methods to the student model, SocialEgoMobile, when 30% of joints and frames are corrupted.
-Distillation based on intermediate features outperforms soft label distillation, as the student model in our experiments must learn to integrate information of additional modalities from high-level social representation provided by the teacher model and compensate for the impact of missing or noisy inputs.
+In Table 2, we compare the performance improvements brought by different knowledge distillation methods to SocialEgoMobile.
+Distillation based on intermediate features outperforms soft label distillation, as the student model in our experiments must learn to integrate information of additional modalities from the teacher’s social representation and compensate for the impact of noisy inputs.
 Learning only from the teacher’s output distributions is insufficient.
-Among the intermediate feature-based methods, InfoNCE (Oord et al., 2018) achieves the best performance.
-By maximizing mutual information between the teacher and student model’s intermediate social representations, InfoNCE encourages the student to learn semantically aligned features, rather than enforcing exact numerical matching at lower levels.
-This is suitable for our experiment setting, where the teacher and student operate on different input modalities and the student receives incomplete and noisy inputs, making low-level alignment impractical.
+Among the feature-based methods, InfoNCE (Oord et al., 2018) achieves the best performance.
+By maximizing mutual information between the teacher and student model’s social representations, InfoNCE encourages the student to learn semantically aligned features, rather than enforcing exact numerical matching at lower levels.
+This is suitable for our experiment setting, where the teacher and student operate on different input modalities and the student receives noisy inputs, making low-level alignment impractical.
 InfoNCE promotes the alignment of high-level semantic information, leading to better noise tolerance and generalization.
-Model Architecture JPL-Social
-Intent F1 Attitude F1 Action F1
-SocialEgoMobile 88.10 80.95 75.00
-w/ 3 layers GAT 84.52 76.19 69.05
-w/ 1 layers GAT 83.33 69.05 67.86
-w/ 2 layer GCN (Kipf and Welling, 2017) 85.71 71.43 67.86
-w/ 2 layers Bi-LSTM 88.10 82.14 75.00
-w/ 1 layer TCN (Lea et al., 2017) 83.33 69.05 65.48
-w/ 1 layer Transformer (Vaswani et al., 2017) 77.38 66.67 54.76
-Table 3.
-Performance Comparison of SocialEgoMobile with Different Backbone Models on JPL-Social when 30% joints and frames are corrupted.
-The best results are shown in bold and the second best results are underlined.
+Table 3. Comparison of SocialEgoMobile with different architectures when 30% joints and frames are corrupted.
+Model Architecture JPL-Social HARPER
+Intent Acc. Attitude Acc. Action Acc. Intent Acc. Attitude Acc. Action Acc.
+SocialEgoMobile 88.10 80.95 75.00 67.31 52.14 57.69
+w/ 1 layers GAT 83.33 69.05 67.86 63.46 43.49 48.09
+w/ 3 layers GAT 84.52 76.19 69.05 65.38 47.83 53.85
+w/ 2 layer GCN (Kipf and Welling, 2017) 85.71 71.43 67.86 59.62 47.83 51.92
+w/ 2 layers Bi-LSTM 88.10 82.14 75.00 61.84 52.14 51.92
+w/ 1 layer TCN (Lea et al., 2017) 83.33 69.05 65.48 59.62 39.13 48.08
+w/ 1 layer Transformer (Vaswani et al., 2017) 77.38 66.67 54.76 55.77 26.08 42.31
 Student Model Architecture.
-To enhance the robustness and efficiency of processing noisy pose feature sequences, we designed the student model SocialEgoMobile with a two-layer GAT (Veličković et al., 2018) followed by a one-layer Bi-LSTM (Graves and Schmidhuber, 2005).
+To enhance the robustness and efficiency of processing noisy pose feature sequences, we designed SocialEgoMobile with a two-layer GAT (Veličković et al., 2018) followed by a one-layer Bi-LSTM (Graves and Schmidhuber, 2005).
 Table 3 presents a comparison of different model configurations under the knowledge distillation with corrupted inputs.
 In the spatial dimension, a two-layer GAT achieves a good trade-off between expressiveness and stability.
-Fewer layers limit the receptive field, reducing the model’s ability to capture global graph structural patterns, while deeper GATs tend to propagate noise more widely.
-Moreover, GAT’s adaptive attention mechanism enables selective aggregation of useful information from neighbouring nodes, offering greater resilience to noisy or missing joints compared to standard GCNs (Kipf and Welling, 2017).
-In the temporal dimension, a single-layer Bi-LSTM provides sufficient modelling capacity for short pose feature sequences, while deeper Bi-LSTMs introduce additional parameters with marginal performance improvement.
-Compared to TCNs (Lea et al., 2017) and Transformers (Vaswani et al., 2017), Bi-LSTM is better suited for sequential data with inherent temporal dependencies and benefits from its gating mechanism, which improves robustness to sparsity and temporal discontinuity.
+Fewer layers limit the receptive field, reducing the ability to capture global graph patterns, while deeper GATs tend to propagate noise more widely.
+Moreover, GAT’s adaptive attention mechanism enables selective aggregation of useful information from neighbouring joints, offering greater resilience to noisy signals compared to standard Graph Convolutional Networks (GCN) (Kipf and Welling, 2017).
+In the temporal dimension, a single-layer Bi-LSTM provides sufficient modelling capacity for short feature sequences, while deeper Bi-LSTMs do not result in a steady performance improvement.
+Compared to Temporal Convolutional Networks (TCN) (Lea et al., 2017) and Transformers (Vaswani et al., 2017), Bi-LSTM is better suited for sequential data with inherent temporal dependencies and benefits from its gating mechanism, which improves robustness to sparsity and temporal discontinuity.
 Overall, the combination of two GAT layers and one Bi-LSTM in SocialEgoMobile offers an effective balance between expressiveness, robustness, and computational efficiency.
@@
 To summarize our results, we highlight three main observations:
-  • The proposed teacher model SocialC3D effectively integrates multimodal inputs (body, face, hand, gaze and raw image) for egocentric social understanding, surpassing SOTAs.
-Each modality contributes unique and complementary information.
-  • The knowledge distillation framework enables the student model to absorb multimodal and clean knowledge from the teacher, improving its ability of capture meaningful information and robustness against spatial-temporal noisy data.
-  • The distilled student SocialEgoMobile achieves a trade-off between performance, robustness, and computational cost, which is suitable for deployment on resource-constrained systems.
+  • SocialC3D surpasses SOTAs by effectively integrating body, face, hand, gaze, and image inputs for egocentric social understanding, demonstrating the value of complementary multimodal cues and the need for socially aware robots capable of comprehensively interpreting human behaviour.
+  • The knowledge distillation framework enables the student model to absorb multimodal knowledge from the teacher, improving its ability to capture meaningful information and robustness against spatial-temporal noisy data in a deployment setting.
+  • The distilled SocialEgoMobile achieves a great trade-off between performance, robustness, and computational efficiency, which is suitable for deployment on resource-constrained systems, promoting practical and scalable HRI in real-world settings.
 5. Conclusion
-In this paper, we address the challenges of performance, robustness, and computational efficiency in forecasting users’ social interaction intentions and behaviours from an egocentric perspective on mobile platforms or embedding system.
-We propose SocialEgoMobile, a lightweight social cue understanding model based on knowledge distillation.
-In our knowledge distillation framework, the teacher model, SocialC3D, uses multimodal inputs, including body poses, raw images, facial expressions, hand gestures, and gaze patterns, while the student model only relies on deliberate corrupted body pose features.
-By distilling knowledge from the teacher model’s integrated multimodal representations and downstream task supervision, the student model learns to produce robust predictions under noisy and incomplete input conditions.
-This work offers an effective and lightweight solution for real-time social understanding, particularly suited for resource-constrained scenarios such as social robots and intelligence social agents.
+In this paper, we address the challenges of performance, robustness, and computational efficiency in forecasting users’ social interaction intentions and behaviours from an egocentric perspective on mobile platforms or embedding systems.
+We propose SocialEgoMobile, a lightweight social understanding model based on knowledge distillation.
+In our knowledge distillation framework, the teacher model, SocialC3D uses multimodal inputs, including body poses, raw images, facial expressions, hand gestures, and gaze patterns, while the student model, SocialEgoMobile only relies on deliberately corrupted body pose features.
+By distilling knowledge from the teacher model’s integrated multimodal social representations and downstream task supervision, the student model learns to produce robust predictions under noisy and incomplete input conditions.
+This work offers an effective and lightweight solution for real-time social understanding in noisy environments, particularly suited for resource-constrained platforms such as social robots and intelligence social agents.
 Future work will focus on further enhancing the model’s robustness through real-world deployment and investigating its generalization to complex scenarios, including domain adaptation across viewpoints and environments.
-This paper has been submitted to ACM Multimedia 2025.